# Revisiting Reactor Anti-Neutrino 5 MeV Bump with 13C Neutral-Current Interaction

### v1

@@ Abstract
For the first time, we systematically investigate the potential of neutrino-nucleus neutral current interactions with 13C to identify the origin of the 5 MeV bump observed in reactor anti-neutrino spectra in the inverse beta decay process. The distinctive signal is obtained from the de-excitation of 13C∗ into the ground state emitting a 3.685 MeV photon in various liquid scintillator detectors. Such an interaction predominantly occurs for the reactor anti-neutrinos within the energy range coinciding with the 5 MeV bump. For a detector that has a capability of 95% level photon and electron separation and small thorium contamination below gr/gr located in a site with an overburden of about a few hundred m.w.e, such as the location of near detectors of RENO and Daya Bay will have a great sensitivity to resolve the 5 MeV bump. In addition, we propose a novel approach to track the time evolution of reactor isotopes by analyzing our 13C signal shedding light on the contributions from 235U or 239Pu to the observed bump. This provides an extra powerful tool in both discriminating the flux models and testing any new physics possibilities for the 5 MeV bump at 3 to 5 level with much less systematic uncertainties and assuming 10 kt.year of data collection. Our detector requirements are realistic, aligning well with recent studies conducted for existing or forthcoming experiments.
Introduction.
@@
Since its first discovery Cowan et al. (1956), reactor neutrinos have advanced our comprehension of the lepton sector. Notably, the KamLAND experiment confirmed the neutrino oscillation as the explanation for the solar neutrino problem Gando et al. (2013). Daya Bay Adey et al. (2018), RENO Bak et al. (2018), and Double Chooz Abe et al. (2014) experiments measured a non-zero value of oscillation parameter . This discovery has opened the possibility of CP violation in neutrino oscillation. Future reactor neutrino experiment JUNO An et al. (2016a) aims to determine the neutrino mass ordering and achieve sub-percent precision in measuring solar neutrino oscillation parameters, which can be complemented by the liquid scintillator counter at Korean new underground lab, Yemilab Seo et al. (2023); Bakhti et al. (2023), promising precise investigations into solar characteristics.
@@
Probes of neutrino oscillation using reactor neutrinos necessitate accurate theoretical predictions of the neutrino flux. However, its calculation is very complicated and traces its origins to the early days subsequent to the first detection of neutrinos in the Cowan and Reines reactor experiment Cowan et al. (1956). The Vogel model Vogel and Engel (1989), employed as the standard flux model for two decades starting from the 1990s, relies on the conversion method and predicts reactor flux measurements based on the Institut Laue-Langevin (ILL) electron spectrum measurements Schreckenbach et al. (1981); Von Feilitzsch et al. (1982).
@@
Interestingly, the actual neutrino fluxes measured in several short-baseline reactor neutrino experiments with varying fission fractions were smaller than the expected values, which has drawn more careful calculations, such as those by Mueller et al. Mueller et al. (2011) and Huber Huber (2011). Nonetheless, the predicted fluxes still surpass the observations by about 6% corresponding to a level discrepancy in the overall energy spectra, now dubbed as the reactor antineutrino anomaly (RAA) Mention et al. (2011).
@@
Moving the focus to a narrower prompt energy range of 4 to 6 MeV, the situation becomes more arduous. In contrast to the deficit in the overall flux, the observed data first reported by RENO Choi et al. (2016), and then confirmed by other experiments such as Daya Bay An et al. (2016b), Double Chooz Abe et al. (2014), NEOS Ko et al. (2017), Neutrino-4 Serebrov et al. (2021), and DANSS (preliminary) Danilov (2022) at more than 4. Moreover, STEREO Almazán et al. (2021) and PROSPECT Andriamirado et al. (2023), utilizing research reactors powered by 100% 235U fuel, have excluded the no 5 MeV bump scenario with more than 3.5 and 2 C.L. respectively. This anomaly is called the 5 MeV bump Huber (2017).
@@
Reactor electron anti-neutrinos are predominantly detected through the inverse beta decay (IBD) reaction, offering distinct advantages such as a large cross-section and a clear coincidence of the prompt positron annihilation and the delayed neutron capture -ray emission. The detection threshold for IBD is 1.8 MeV. The previously mentioned 5 MeV bump is observed from IBD, which means the positron energies peak around 5 MeV beyond the expectation by Huber-Mueller (HM). To further investigate its origin, whether it is coming from a miscalculation of the flux, new physics, or IBD systematic, employing another detection approach is advantageous. In this letter, for the first time, we propose to use the neutral current (NC) interaction of neutrinos with 13C isotope ubiquitous with mostly about 1.1 % abundance in the carbon-based liquid scintillator (LS) detectors to identify the 5 MeV bump alternatively to the IBD. The interaction threshold is 3.685 MeV, resulting in the excitation of 13C nuclei to their first excited state. Then, a 3.685 MeV -ray emits from prompt de-excitation of 13C∗ nuclei, creating a distinctive signal. Note that NC interactions are flavor blind and are not affected by neutrino oscillation parameter uncertainties. However, other types of interactions such as neutrino-electron elastic scattering (ES) and coherent elastic neutrino-nucleus scattering (CENS), which benefit from low energy thresholds, suffer from large systematic uncertainty, and detecting signals above the background poses significant challenges rendering them less suitable for studying the 5 MeV bump. It is worth noting that CC and NC interactions of deuterons can be alternative methods in experiments with a sufficient quantity of heavy water 111A review of detection channels for reactor antineutrinos is found in Ref. Qian and Peng (2019)..
Expected event spectrum.
@@
The cross-section for the NC interaction of neutrinos with 13C is given by Suzuki et al. (2012):
@@
where MeV is the energy threshold, , and are constants Fukugita et al. (1988); Suzuki et al. (2019). Figure 1 shows the cross-section represented by the dashed green curve. Notice that for light nuclei such as 13C, there is a good agreement between theory and experiment. Hence, we expect the cross-section uncertainty would follow the best theoretical calculations to the 1% level Barrett et al. (2013); Zhao et al. (2022) although its experimental measurement lacks so far. Various experiments, including reactor neutrino interactions, solar neutrino observations at JUNO Zhao et al. (2022) and Yemilab Seo et al. (2023); Bakhti et al. (2023), and electron anti-neutrino production at IsoDAR Alonso et al. (2022); Seo et al. (2023) can measure the cross-section in the near future Bakhti et al. .
@@
In Fig. 1, we present the flux of reactor neutrinos with the dashed blue curve, assuming the HM model Huber (2011); Mueller et al. (2011) and considering the average fission fractions at near detector of RENO Yoon et al. (2021) as follows: 0.571 for 235U, 0.073 for 238U, 0.3 for 239Pu, and 0.056 for 241Pu. Other reactor models are not illustrated as they overlap with each other. Moreover, we show the product of the flux and the interaction cross-section with 13C, considering various flux models such as HM, Vogel Vogel and Engel (1989), KI (Kurchatov Institute) Kopeikin et al. (2021), EF (Estienne-Fallot) Estienne et al. (2019), and HKSS (Hayen-Kostensalo-Severijns-Suhonen) Hayen et al. (2019), as well as the reconstructed number of events considering excess events at RENO Choi et al. (2016). For a review of different models, see Giunti et al. (2022); Zhang et al. (2024).
@@
As demonstrated in Fig. 1, a significant number of events falls within the neutrino energy range of 4.5 to 7.5 MeV, coinciding with the 5 MeV bump anomaly. This interaction thus presents a novel opportunity to investigate the differentiate among reactor flux models, offering a new channel testing the phenomenon. The ratios of events for the Vogel, KI, EF, and HKSS models to HM model are 0.96, 0.96, 0.97, and 1.03, respectively 222Note that the HKSS model is introduced to fit the 5 MeV bump.. Assuming all the excess events in the IBD measurements as reported in Ref. Zhang et al. (2024) are caused by a higher flux, the experimental events involving will likely be 4 to 8 higher than predicted by the HM model. In the case of RENO Yoon et al. (2021), as demonstrated in Fig. 1, it is higher than the HM model events. Hence investigations identifying the origin of the 5 MeV bump rely on how much we can reduce the uncertainties in the current/future experiments.

### v2

@@ Abstract
For the first time, we comprehensively examine the potential of neutrino-nucleus neutral current interactions with 13C to determine the origin of the 5-MeV bump in reactor antineutrino spectra observed through the inverse beta decay (IBD) process. This anomaly may be due to new physics, reactor antineutrino flux inaccuracies, or IBD systematics. The primary signal is the 3.685 MeV photon released during the de-excitation of 13C∗ to its ground state, detectable in various liquid scintillator detectors. This interaction is most common for reactor antineutrinos within the energy range of the 5-MeV bump, providing unique sensitivity to its origins. Remarkably, we confirm the powerfulness of our proposal by completely ruling out a new physics scenario explaining the bump from the existing NEOS data. The main systematic challenge in distinguishing flux models is the accuracy of cross-section measurements; thus, minimizing these uncertainties is essential. Hence, we explore the potential of current and forthcoming experiments, including solar neutrino studies at JUNO, pion and muon decay-at-rest experiments at OscSNS, and isotope decay-at-rest studies at Yemilab, to obtain precise cross-section measurements. Additionally, we propose a novel method to track the time evolution of reactor isotopes by analyzing the 13C signal. This technique yields critical insights into the contributions of 235U and 239Pu to the observed bump, acting as a robust tool to differentiate between flux models and explore new physics explanations for the 5-MeV bump.
I Introduction.
@@
Since its first discovery [1], reactor neutrinos have advanced our comprehension of the lepton sector. Notably, the KamLAND experiment confirmed the neutrino oscillation as the explanation for the solar neutrino problem [2]. Daya Bay [3], RENO [4], and Double Chooz [5] experiments measured a non-zero value of oscillation parameter . This discovery has opened the possibility of CP violation in neutrino oscillation. Future reactor neutrino experiment JUNO [6] aims to determine the neutrino mass ordering and achieve sub-percent precision in measuring solar neutrino oscillation parameters, which can be complemented by the liquid scintillator counter at Korean new underground lab, Yemilab [7, 8], promising precise investigations into solar characteristics.
@@
Probes of neutrino oscillation using reactor neutrinos necessitate accurate theoretical predictions of the neutrino flux. However, its calculation is very complicated and traces its origins to the early days subsequent to the first detection of neutrinos in the Cowan and Reines reactor experiment [1]. The Vogel model [9], employed as the standard flux model for two decades starting from the 1990s, relies on the conversion method and predicts reactor flux measurements based on the Institut Laue-Langevin (ILL) electron spectrum measurements [10, 11].
@@
Interestingly, the actual neutrino fluxes measured in several short-baseline reactor neutrino experiments with varying fission fractions were smaller than the expected values, which has drawn more careful calculations, such as those by Mueller et al. [12] and Huber [13]. Nonetheless, the predicted fluxes still surpass the observations by about 6% corresponding to a level discrepancy in the overall energy spectra, now dubbed as the reactor antineutrino anomaly (RAA) [14].
@@
Moving the focus to a narrower prompt energy range of 4 to 6 MeV, the situation becomes more arduous. In contrast to the deficit in the overall flux, the observed data first reported by RENO [15], and then confirmed by other experiments such as Daya Bay [16], Double Chooz [5], NEOS [17], Neutrino-4 [18], and DANSS (preliminary) [19] at more than 4. Moreover, STEREO [20] and PROSPECT [21], utilizing research reactors powered by 100% 235U fuel, have excluded the no 5 MeV bump scenario with more than 3.5 and 2 C.L. respectively. This anomaly is called the 5 MeV bump [22].
@@
Reactor electron anti-neutrinos are predominantly detected through the inverse beta decay (IBD) reaction, offering distinct advantages such as a large cross-section and a clear coincidence of the prompt positron annihilation and the delayed neutron capture -ray emission. The detection threshold for IBD is 1.8 MeV. The previously mentioned 5 MeV bump is observed from IBD, which means the positron energies peak around 5 MeV beyond all the theoretical expectation including the Huber-Mueller (HM) model. To further investigate its origin, whether it is coming from a miscalculation of the flux, new physics [23], or IBD systematic, employing another detection approach is advantageous. In this paper, for the first time, we propose to use the neutral current (NC) interaction of neutrinos with 13C isotope ubiquitous with mostly about 1.1 % abundance in the carbon-based liquid scintillator (LS) detectors to identify the 5 MeV bump alternatively to the IBD. The interaction threshold is 3.685 MeV, resulting in the excitation of 13C nuclei to their first excited state. Then, a 3.685 MeV -ray emits from prompt de-excitation of 13C∗ nuclei, creating a distinctive signal. Note that NC interactions are flavor blind and are not affected by neutrino oscillation parameter uncertainties. However, other types of interactions such as neutrino-electron elastic scattering (ES) and coherent elastic neutrino-nucleus scattering (CENS), which benefit from low energy thresholds, suffer from large systematic uncertainty, and detecting signals above the background poses significant challenges rendering them less suitable for studying the 5 MeV bump. It is worth noting that CC and NC interactions of deuterons can be alternative methods in experiments with a sufficient quantity of heavy water 111A review of detection channels for reactor antineutrinos is found in Ref. [25]..
The origin of the observed 5-MeV spectral bump may be attributed to new physics, inaccuracies in IBD systematics, or errors in flux calculations. In the context of new physics, the observed 5-MeV bump could be explained either by modifications to the flux — such as the production of new particles that decay into —or through mechanisms that do not alter the flux. In the first case, distinguishing this scenario from a standard model (SM) flux miscalculation using 13C events is not feasible. However, in the second case, such mechanisms would be distinguishable through the 13C signal. In this regard, we validate the effectiveness of our alternative channel by directly searching for the 3.685 MeV photon signal in the recent NEOS results, conclusively ruling out the new physics explanation of the 5-MeV bump such as the scenario involving a sterile neutrino with an additional U(1) gauge symmetry [23]. This conclusion arises from the fact that such a new physics scenario would predict a significant, sharp photon peak around 3.685 MeV, well above the background events, resulting from sterile neutrino scattering with 13C mediated by a light boson. The absence of this peak in the existing data strongly disfavors this explanation.
In case that bump results from IBD systematics, measurements of the 13C flux are expected to align with flux models. As an example of IBD systematics, reconstructing prompt energy in liquid scintillators involves several complexities. The light yield demonstrates non-linearity with respect to the energy deposited by charged particles, primarily due to the quenching effect, as described by Birks’ law [25, 26], and contributions from Cherenkov radiation. Furthermore, electronic readout systems may introduce additional non-linearity by inaccurately estimating the number of detected photoelectrons. Variations in detector response, depending on the event position, further exacerbate these challenges. Consequently, comprehensive calibration procedures are critical for accurate prompt energy reconstruction. To address the 5-MeV bump effectively and verify that it is not a systematic artifact, alternative methods, such as analyzing 13C events, are indispensable.
To assess the broader feasibility of our proposed channel beyond the specific scenario discussed, we perform a detailed analysis of background events and calculate the signal-to-background ratio required to differentiate between various reactor neutrino flux models. This analysis is presented in the appendix. Moreover, the most significant systematic uncertainty in differentiating reactor flux models arises from cross-section measurements. To address this issue, we explore the potential of current and future experiments—such as solar neutrino studies at JUNO, pion and muon decay-at-rest experiments at OscSNS, and isotope decay-at-rest studies at Yemilab—to achieve precise and accurate cross-section measurements. Additionally, we assess the feasibility of measuring the cross section under various reactor models to enhance our understanding and improve the reliability of flux model separation.
The structure of the paper is as follows. In Sec. II, we discuss the expected number of events for 13C. Section III examines a Beyond the Standard Model (BSM) scenario that addresses the 5-MeV bump using 13C events at NEOS. In Sec. IV, we provide a detailed analysis of the 13C cross-section measurement. Section V investigates the potential for differentiating reactor flux models and their application in detecting fission fraction evolution. Finally, our conclusions are presented in Sec. VI. The detailed explanations on the background analysis along with an experimental proposal realizing our goal are summarized in Appendix A .
II Cross section of 13C NC and Expected events.
In this section, we discuss the Cross section and expected signal events associated with 13C in carbon-based liquid scintillators.
@@
The cross section for the neutral current interaction of neutrinos with 13C is given by [27]:
@@
where MeV is the energy threshold, , and are constants [28, 29]. Fig. 1 shows the cross section represented by the dashed green curve. Notice that for light nuclei such as 13C, there is a good agreement between theory and experiment. Hence, we display Fig. 1 assuming that the cross section uncertainty would follow the best theoretical calculations to the 1% level [30, 31] although its experimental measurement lacks so far. Following the solar neutrino flux observation proposals using the 13C NC interaction [32, 33, 31], we expect to measure the cross section within a few % level uncertainty in the near future as explained in Sec. IV.
@@
In Fig. 1, we present the flux of reactor neutrinos with the dashed blue curve, assuming the HM model [13, 12] and considering the average fission fractions at near detector of RENO [37] as follows: 0.571 for 235U, 0.073 for 238U, 0.3 for 239Pu, and 0.056 for 241Pu. Other reactor models are not illustrated as they overlap with each other. Moreover, we show the product of the flux and the interaction cross-section with 13C, considering various flux models such as HM, Vogel [9], KI (Kurchatov Institute) [36], EF (Estienne-Fallot) [35], and HKSS (Hayen-Kostensalo-Severijns-Suhonen) [34], as well as the reconstructed number of events considering excess events at RENO [15]. For a review of different models, see [38, 39].
@@
As demonstrated in Fig. 1, a significant number of events falls within the neutrino energy range of 4.5 to 7.5 MeV, coinciding with the 5 MeV bump anomaly. This interaction thus presents a novel opportunity to investigate the origin of the bump and possibly differentiate among reactor flux models, offering a new channel testing the phenomenon. The ratios of events for the Vogel, KI, EF, and HKSS models to HM model are 0.96, 0.96, 0.97, and 1.03, respectively 222Note that the HKSS model is introduced to fit the 5 MeV bump.. Assuming all the excess events in the IBD measurements as reported in Ref. [39] are caused by a higher flux, the experimental events involving 13C will likely be 4 to 8 higher than predicted by the HM model. In the case of RENO [37], as demonstrated in Fig. 1, it is 12% higher than the EF or KI models which fit the RAA within about 1 level. Hence investigations identifying the origin of the 5 MeV bump rely on how much we can reduce the uncertainties in the current/future experiments.